# Neural Video Fields Editing

### 12 Dec 2023 v1

@@ Abstract
Diffusion models have revolutionized text-driven video editing. However, applying these methods to real-world editing encounters two significant challenges: (1) the rapid increase in graphics memory demand as the number of frames grows, and (2) the inter-frame inconsistency in edited videos. To this end, we propose NVEdit, a novel text-driven video editing framework designed to mitigate memory overhead and improve consistent editing for real-world long videos. Specifically, we construct a neural video field, powered by tri-plane and sparse grid, to enable encoding long videos with hundreds of frames in a memory-efficient manner. Next, we update the video field through off-the-shelf Text-to-Image (T2I) models to impart text-driven editing effects. A progressive optimization strategy is developed to preserve original temporal priors. Importantly, both the neural video field and T2I model are adaptable and replaceable, thus inspiring future research. Experiments demonstrate that our approach successfully edits hundreds of frames with impressive inter-frame consistency.
Figure 1: NVEdit enables various editing options, including shape variation, scene change, and style transfer, while preserving the motion and semantic layout of the original scene. Due to its efficient encoding rate, long videos with hundreds of frames even be edited well.
@@ 1 Introduction
Text-driven video editing aims to faithfully edit high-quality videos following the user-given prompts. Due to the limited paired video-prompt data, using videos to train models is expensive [39, 33]. Considering the advancement of Text-to-Image (T2I) methods [27, 11, 22], adopting the pre-trained T2I models to edit videos has attracted wide attention [8, 26, 46, 5]. However, T2I models inherently lack cross-frame semantic consistency, resulting in inevitable flickering between edited frames. Hence, editing videos more temporal-consistently has become a challenging task.
@@
In pursuit of this objective, numerous algorithms improve cross-attention operations [47, 49] and DDIM [41] inversion [25, 4, 38] to ensure superior generation and inter-frame coherence. However, these methods still exhibit two principal deficiencies. First, although various constraints and attention control are introduced to the generation process, inter-frame inconsistency still exhibits in the edited videos, such as object deformation and texture variation. That is because T2I models inherently lack temporal motion priors. The second drawback is the rapid increase in graphics memory demands as more frames are edited, while limited hardware devices constrain the length of videos they can edit. Since editing long videos with hundreds of frames is significant for real-world applications, existing T2I models cannot achieve satisfactory effects in videos. Neural representation [24, 7, 19] may be a feasible tool to decrease memory overhead while improving inter-frame consistency, since it has a high compression rate for videos and models contents well. Some current works have tried this proposal [15, 29] but brought other limitations, such as inaccuracies in complex motion or viewpoint changing, etc.
@@
To address these challenges, a memory-efficient framework is developed for consistent text-driven video editing, avoiding intricate attention-related operations or fine-tuning for T2I models. Inspired by [10], validating the use of a T2I model for updating neural radiance fields in editing 3D scenes, we extend this concept to video signals through a Neural Video Field (NVF). Specifically, our approach comprises two stages: video fitting and field editing. In the first stage, we construct an NVF with tri-plane encoding and sparse grids, efficiently modeling temporal and content priors of a given video. Benefiting from its encoding efficiency, even lengthy videos with hundreds of frames can be represented as a signal field. In the second stage, the trained NVF is updated to incorporate text-based editing effects. During each iteration, NVF renders a frame, and a T2I model is employed to edit it based on the provided text. These edited frames serve as pseudo Ground Truths (GTs) to optimize the parameters of NVF. Since NVF has been initially trained on the original video, it retains robust temporal priors post-optimization while manifesting the desired editing effects. This strategy of using Neural Video fields for video Editing is termed NVEdit. Our NVEdit possesses inherent intriguing properties of neural representation, e.g., frame interpolation, which has been explored in recent works [7, 17]. During inference, NVEdit predicts frames not present in the original video seamlessly, while exhibiting coherent editing effects. In addition, both the NVF and the T2I model within NVEdit are adaptable, allowing for seamless improvements or replacements based on diverse requirements. In this paper, we majorly adopt Instruct-Pix2Pix (IP2P) [3] for editing effects. We observed that for local editing, IP2P tends to modify regions that should be preserved inaccurately, impacting video editing outcomes. To address this, an auxiliary mask is designed to constrain the editable region without additional calculation, resulting in Instruct-Pix2Pix+ (IP2P+). This version not only enhances local editing capabilities but also demonstrates the potential for T2I model improvement within NVEdit. Overall, our contributions are as follows:
@@ 2.1 Text-driven Video Editing
Editing videos through user-given texts has attracted wide attention. A straightforward idea is to train a model with video-text pairs. Ho et al. [13] first proposed to train a diffusion model from videos and images but required paired video-text data, which is expensive to obtain. To reduce the reliance on video data, Singer et al. [39] learned visual-text correlation from paired image-text data and captured motion information from videos. Qin et al. [33] synthesized video-prompt triplets with off-the-shelf models, but training with video data is expensive. As the advancement of conditional diffusion models [35, 1, 30] promotes text-to-image generation [18, 36, 42, 28, 52, 37], exploiting their effective vision priors to edit videos has attracted wide attention. The key challenge is ensuring inter-frame consistency. Wu et al. [47] fine-tuned a T2I model to fit a specific video for temporal modeling. Khachatryan et al. [16] simulated motions by translating latent. Qi et al. [32] fused attention features before and after editing for consistency. Wang et al. [43] developed dense spatial-temporal attention to learn motion. Geyer et al. [9] only edited key frames and propagated to the whole video. Liew et al. [20] decoupled content, structure, and motion signals to model coherent features. However, these methods require rapid increased memory to edit more frames, limiting the potential for editing long videos.
@@ 2.2 Neural Representation for Videos
Recently, some attempts at video editing appeared. Kasten et al. [15] utilized Multilayer Perceptrons (MLPs) to map video pixels to Atlas maps, and edited videos by editing the Atlas maps. Further, Ouyang et al. [29] mapped video content to a canonical image and edited it to realize video editing. However, the strategy of extracting video content to a representation medium and editing it naturally has two disadvantages. First, for videos with significant content changes (such as large movements or long range motion), it is hard to map all content to a single medium, which limits accuracy. Second, their editing effect relies entirely on the image editing of the medium. But T2I models may perform poorly in some cases, i.e., lack robustness.
@@ 3.1 Preliminaries
$\underset{\boldsymbol{\theta}}{\min}\,\mathbbm{E}_{\textbf{z}_{0},\boldsymbol{\varepsilon}\sim\mathcal{N}(0,I),t}||\boldsymbol{\varepsilon}-\varepsilon_{\boldsymbol{\theta}}(\textbf{z}_{t},t,\textbf{c}_{P})||_{2}^{2}.\end{aligned}$ (1)
During inference, given a random latent $\textbf{z}_{T}$ and a prompt P, the trained model predicts noise $\varepsilon_{\boldsymbol{\theta}}(\cdot)$ for $\it{T}$ steps to generate an image, which meets the textual condition P.
@@
where $s_{I}$ and $s_{P}$ are the weights for image and text conditions respectively. The first term means random sampling, the second term guides sampling based on the image condition, while the third term introduces the textual condition.
@@ 3.2 NVEdit
Model structure. NVP [17] validates that the explicit-implicit hybrid structures benefit fitting speed and accuracy. We thus build a hybrid model to represent video signals. As shown in the left part of Fig. 2, our explicit module contains tri-plane and sparse grids. The former is three feature maps designed to record video features and the latter is a set of parameter grids sparsely distributed across the $x-y-t$ coordinate system of the video. Both of them are trainable parameters, denoted as $\boldsymbol{\theta}_{ex}$. Given a spatio-temporal coordinate $(x,y,t)$, we extract features from its neighbor grids and tri-plane, represented as the white cube in Fig. 2. Our implicit part is an MLP that decodes the aggregated features into a pixel value, whose parameters are $\boldsymbol{\theta}_{im}$. We update $\widetilde{\boldsymbol{\theta}}=\left\{\boldsymbol{\theta}_{im},\boldsymbol{\theta}_{ex}\right\}$ to fit a video V and achieve editing effects.
@@
As the batch size is independent of the video length, our graphics memory demand remains constant. Notably, we supervise training with randomly sampled pixels across the entire video rather than a batch of frames, since we found this scheme yields faster convergence and slightly better fitting accuracy compared to training frame-by-frame. In this stage, NVF learns the temporal and motion priors of the given video, which facilitates subsequent consistent editing.
@@
Note that we only update NVF and fix the T2I model. Different from Sec. 3.2 that samples pixels randomly, here we supervise NVF frame-by-frame because T2I models exhibit optimal performance on complete images rather than irregularly sampled pixels. As only one frame is edited per iteration, our graphics memory requirement is constant and independent of the video length. Importantly, we adopt IP2P [3] as the T2I model, and further improve it for better editing effects, which is discussed in Sec. 3.3. It also proves that the T2I model of NVEdit is replaceable and improvable, related further experiments are given in Sec. 4.4.
In addition, a progressive optimization strategy is developed for temporal coherence. Specifically, At the beginning of the field editing stage, we assign a lower weight to the instruction condition, i.e., setting $s_{P}$ in Eq. 2 to a small number, and gradually increasing it as the optimization proceeds. This progressive strategy systematically guides the rendered frames toward the desired editing effects step-by-step, diminishing the updating intensity to NVF at each iteration. Thus, we effectively preserve the temporal priors of the NVF learned at the video fitting stage. Experiments validate the superior temporal consistency of our method.
@@
where T is the set of all coordinates on the temporal axis.
@@
3.3 Instruct-Pix2Pix+
Although IP2P enables impressive image editing effects, for local editing, it alters the region to be retained, as shown in Fig. 3. We suppose this is an inherent flaw of IP2P, as it samples based on Eq. 2, which only ensures results to be roughly consistent with the image condition, but lacks strong constraints on the unedited region. Typically, users prefer instructing specific regions while leaving others untouched. To this end, we develop Instruct-Pix2Pix+ to enhance the content fidelity of edited images to the conditional image $I$, without any additional computation.
@@ 3.3 Instruct-Pix2Pix+
We observed that $\boldsymbol{\varepsilon}_{{\boldsymbol{\theta}}_{P}}$ represents the relationship between the generated image and conditional text. Hence, a simple yet effective auxiliary mask is proposed. We apply normalization and dimensionality reduction to $\boldsymbol{\varepsilon}_{{\boldsymbol{\theta}}_{P}}$, getting a gray-scale image $\textbf{G}_{x,I,T}$, in which a higher value $x$ means the pixel is more likely to be relevant to instruction, and we allow it to change. In contrast, a lower $x$ signals that the pixel is unlikely to require change. We set a mask threshold $\tau\in[0,1]$ on $\textbf{G}_{x,I,T}$ to get our auxiliary mask, $\textbf{M}_{x,I,T}=\mathbbm{1}(\textbf{G}_{x,I,T}\geq\tau)$. Notice that we only use $\boldsymbol{\varepsilon}_{{\boldsymbol{\theta}}_{P}}$, which is an essential element predicted in IP2P, to get $\textbf{M}_{x,I,T}$, without any additional calculation.
@@
Table 1: Quantitative evaluation of the 16 sets of edited results obtained from different SOTA approaches proposed in 2023, including IP2P [3], Video-P2P [21], V2V-zero [43], TokenFlow [9], Fatezero [32], and CoDeF [29]. ‘Tem-Con’, ‘Frame-Acc’, and ‘Frame-Score’ are calculated by the pre-trained CLIP [34] model, while ‘Edit’, ‘Image’, and ‘Temporal’ are obtained from users. The best and the second-best scores are highlighted in bold and underline respectively.
4 Experiments
4.1 Implementation Details
To generate edited frames for optimizing NVF, we directly use the pre-trained IP2P [3] as the base model. The proposed auxiliary mask is introduced to improve its local editing capability. Following previous works [17, 10, 32], UVG-HD [23], DAVIS [31], and some other in-the-wild videos are adopted to evaluate the effect of our approach.
@@ 4.2 Comparison with the state-of-the-art
We prove our superior effectiveness by comparing with other five State-Of-The-Art (SOTA) text-driven video editing approaches, including Video-P2P [21], Vid2Vid-zero (V2V-zero) [43], TokenFlow [9], Fatezero [32], and CoDeF [29]. Furthermore, to demonstrate that the proposed method generates more temporal consistent results than directly applying the T2I model on videos, we fix the random seed of IP2P [3] and showcase its results as a baseline. All of these methods are recently proposed in 2023. Note that for prompt-to-prompt approaches, to ensure a fair comparison, we elaborate the source-target prompt pairs for them.
Quantitative comparison. We present both the objective and subjective metrics of our method and other recent SOTA algorithms. For subjective evaluation, we prepared 16 sets of edited results, shuffled the order of methods within each set, and asked users to rate different results in each set from 1 (the worst) to 5 (the best). As shown in Tab. 1, the proposed method achieves the best temporal consistency and frame-wise editing accuracy, and demonstrates comparable semantic alignment to the canonical image-based method [29] in the feature space of the CLIP model. As for user studies, on average, our approach obtains the best user preferences across three aspects.
@@
Figure 5: NVEdit enables temporal consistent, high-quality edits of real-world long videos with hundreds of frames. The edited results not only faithfully conform to the instruction, but also strictly inherit original information, such as the posture, background, and motion.
@@
Figure 6: Comparison of graphics memory overhead. One can see that as frames increase, our method only adds minimal memory overhead, which allows it to edit long videos.
@@
Qualitative comparison. In order to provide a more intuitive comparison, we present the visual results of NVEdit and other baselines in two different scenarios in Fig. 4, showcasing two types of motion changes: the upper scenario changes action, while the lower one changes viewpoint. Instructions used in IP2P and our method are given above. For the prompt-to-prompt methods, we designed original prompts as ‘A man is walking his dog on the road in front of the courtyard’ and ‘A person stands before the wall’, whilst the target prompts are ‘A man is walking his tiger on the road in front of the courtyard’ and ‘A person in pink shirt stands before the wall’. Observing the results, the edited results of frame-wise IP2P vary a lot among different frames (as indicated by the arrows) and fail to preserve unedited region (as shown in the bottom scene). Although other methods develop various techniques, such as attention-related control or null-text inversion, their results incorrectly modify the background and posture, or fail in certain viewpoints. Notice that CoDeF maps video content to a canonical image and edits it through ControlNet [50]. For a fair comparison, we use IP2P to edit the canonical image of CoDeF in this paper, omitting the proposed auxiliary mask as it is a novel tool invented in this work. In the top-row example, the editing of the canonical image is occasionally subpar. However, CoDeF relies entirely on this edit, leading to an unsatisfactory appearance in all of its predicted frames. In contrast, our method, with progressive optimization using numerous pseudo-GTs, accommodates errors in certain pseudo-GTs. Besides, in the lower-row example, IP2P incorrectly colors the wall, consequently impacting the entire video of CoDeF, whereas our IP2P+ precisely controls the editable region to improve video editing.
@@
Graphics memory comparison. To demonstrate the superior graphics memory efficiency of our method, we record the graphics memory overhead of different algorithms using frames with a resolution of 480 $\times$ 480. The experiments are conducted on a single NVIDIA A40 GPU. Due to its limited graphics memory, some methods cannot edit videos with more than 16 frames. Therefore, we selected 4, 8, 12, and 16 frames for our analysis. Results are displayed in a line chart to illustrate the memory growth trend of different methods. As depicted in Fig. 6, except CoDeF and our method, other approaches require inputting more frames to diffusion models for editing, resulting in a rapid increase in graphics memory. Although TokenFlow incurs a smaller cost than ours, its growth trend is evident, limiting its ability to edit long videos. Leveraging the efficient encoding of neural representation, both CoDeF and our method add only minor additional memory. Since CoDeF [29] employs an implicit representation, while we develop an explicit-implicit hybrid structure as discussed in Sec. 3.2, our graphics memory overhead is higher than that of CoDeF. However, as noted in [6], our hybrid model enables a better representation effect than the implicit CoDeF. Besides, CoDeF maps video content to a canonical image, encoding only motion information, which reduces graphics memory demand. But for videos with significant variations, its mapping and encoding tend to be inaccurate. In contrast, our method encodes content and motion together through an effective hybrid structure, addressing this problem.
[…]

### 09 Mar 2024 v2

@@ Abstract
Diffusion models have revolutionized text-driven video editing. However, applying these methods to real-world editing encounters two significant challenges: (1) the rapid increase in GPU memory demand as the number of frames grows, and (2) the inter-frame inconsistency in edited videos. To this end, we propose NVEdit, a novel text-driven video editing framework designed to mitigate memory overhead and improve consistent editing for real-world long videos. Specifically, we construct a neural video field, powered by tri-plane and sparse grid, to enable encoding long videos with hundreds of frames in a memory-efficient manner. Next, we update the video field through off-the-shelf Text-to-Image (T2I) models to impart text-driven editing effects. A progressive optimization strategy is developed to preserve original temporal priors. Importantly, both the neural video field and T2I model are adaptable and replaceable, thus inspiring future research. Experiments demonstrate the ability of our approach to edit hundreds of frames with impressive inter-frame consistency. Our project is available at: https://nvedit.github.io/.
Keywords:
Video Editing Neural Video Field T2I Model
Figure 1: NVEdit enables various editing options, including shape variation, scene change, and style transfer, while preserving original motion and semantic layout. Due to its efficient encoding rate, long videos with hundreds of frames even be edited well.
@@ 1 Introduction
Text-driven video editing aims to faithfully edit high-quality videos following the user-given prompts. Considering the advancement of Text-to-Image (T2I) methods [31, 16, 26], adopting the pre-trained T2I models to edit videos has attracted wide attention [12, 30, 50, 6]. However, T2I models inherently lack temporal priors, resulting in inevitable inconsistency across edited frames. Hence, editing videos more temporal-consistently has become a challenging task.
@@
In pursuit of this objective, numerous algorithms improve cross-attention operations [52, 53] and DDIM [45] inversion [29, 5, 42, 55] to ensure superior generation while maintaining inter-frame coherence. However, these methods still exhibit two principal deficiencies. First, although various constraints and attention controls are introduced to the generation process, inter-frame inconsistency still exhibits in the edited videos, such as object deformation and texture variations. That is because T2I models inherently lack temporal motion priors. The second drawback is the rapid increase in GPU memory demands as more frames are edited, while limited hardware devices constrain the length of videos they can edit. Since editing long videos with hundreds of frames is significant for real-world applications, existing T2I models cannot achieve satisfactory effects in videos. Neural representation [28, 8, 23] may be a feasible tool to decrease memory overhead while improving inter-frame consistency, as it is naturally a smooth representation [44] and has a high compression rate for videos [8]. Some current works have attempted this proposal [20, 33] but brought other limitations, such as inaccuracies in complex motion or viewpoint changing, etc.
@@
To address these challenges, a memory-efficient framework is developed for consistent text-driven video editing, avoiding intricate attention-related operations or fine-tuning for T2I models. Specifically, our approach comprises two stages: video fitting and field editing. In the first stage, we construct a Neural Video Field (NVF) with tri-plane encoding and MultiLayer Perceptron (MLP) decoding, efficiently modeling temporal and content priors of a given video. Benefiting from its encoding efficiency, even lengthy videos with hundreds of frames can be compactly represented as a signal field. In the second stage, the trained NVF is updated to incorporate text-based editing effects. During each iteration, NVF renders a frame, and a T2I model is employed to edit it based on the provided text. These edited frames serve as pseudo Ground Truths (GTs) to optimize the parameters of NVF. Since NVF has been initially trained on the original video, it retains robust temporal priors after optimization while manifesting the desired editing effects. This strategy of using Neural Video fields for video Editing is termed NVEdit. Our NVEdit also possesses inherent intriguing properties of neural representation, e.g., frame interpolation, which has been explored in recent works [8, 22]. It means that NVEdit can predict frames not present in the original video while exhibiting coherent editing effects. In addition, both the NVF and the T2I model within NVEdit are adaptable, allowing for seamless improvements or replacements based on diverse requirements. In this paper, we majorly adopt InstructPix2Pix (IP2P) [4] for editing effects. We observed that for local editing, IP2P tends to modify regions that should be preserved inaccurately, impacting video editing outcomes. To address this, an auxiliary mask is designed to constrain the editable region without additional calculation, resulting in Instruct-Pix2Pix+ (IP2P+). This version not only enhances local editing capabilities but also demonstrates the potential for T2I model improvement within NVEdit. Overall, our contributions are as follows:
@@ 2.1 Text-driven Video Editing
Editing videos through user-given texts has attracted wide attention [18, 46]. A straightforward idea is to train a model with video-text pairs. Ho et al. [18] first proposed to train a diffusion model from videos and images but required paired video-text data, which is expensive to obtain. To reduce the reliance on video data, Singer et al. [43] learned visual-text correlation from paired image-text data and captured motion information from videos. Qin et al. [37] synthesized video-prompt triplets with off-the-shelf models, but training with video data is expensive. As the advancement of conditional diffusion models [39, 1, 34] promotes text-to-image generation [32, 40, 41], exploiting their effective vision priors to edit videos has attracted wide attention. The key challenge is ensuring inter-frame consistency. Wu et al. [52] fine-tuned a T2I model to fit a specific video for temporal modeling. Khachatryan et al. [21] simulated motions by translating latent. Qi et al. [36] fused attention features before and after editing for consistency. Wang et al. [47] developed dense spatial-temporal attention to learn motion. Geyer et al. [13] only edited key frames and propagated to the whole video. Liew et al. [24] decoupled content, structure, and motion signals to model coherent features. However, these methods require rapid increased memory to edit more frames, limiting the potential for editing long videos.
@@ 2.2 Neural Representation for Videos
Recently, some attempts at video editing appeared. Kasten et al. [20] utilized MLPs to map video pixels to Atlas maps, and edited videos by operating on Atlas maps. Further, Ouyang et al. [33] mapped video content to a canonical image and edited it to realize video editing. However, the strategy of extracting video content to a representation medium and editing it naturally has two disadvantages. First, for videos with significant content changes (such as large movements or long range motion), it is hard to map all content to a single medium accurately. Second, their editing effect entirely relies on the image editing of the medium. But T2I models may perform poorly in certain cases, i.e., lacking robustness.
@@ 3.1 Preliminaries
$\underset{\boldsymbol{\theta}}{\min}\,\mathbbm{E}_{\textbf{z}_{0},\boldsymbol{\varepsilon}\sim\mathcal{N}(0,I),t}\|\boldsymbol{\varepsilon}-\varepsilon_{\boldsymbol{\theta}}(\textbf{z}_{t},t,\textbf{c}_{P})\|_{2}^{2}.\end{aligned}$ (1)
During inference, given a random latent $\textbf{z}_{T}$ and a prompt P, the trained model predicts noise $\varepsilon_{\boldsymbol{\theta}}(\cdot)$ for $\it{T}$ steps to generate an image, which meets the textual condition P. This accomplishment has prompted further exploration in image generation [56], image editing [4], video generation [14], and protection [58].
@@
where $s_{I}$ and $s_{P}$ are the weights for image and text conditions respectively. The first term means random sampling, the second term guides sampling based on the image condition, while the third term introduces the textual condition. Compared to previous text-driven methods [16, 11], which mainly belong to the prompt-to-prompt type and need well-designed paired texts, IP2P enables users to edit images more conveniently, i.e., directly using one oral instruction to edit. Existing method [15] has exploited the prior of IP2P to make 3D editing easier. However, there is still no research on directly benefiting from the off-the-shelf IP2P to conveniently edit videos with oral instructions.
@@ 3.2 NVEdit
Model structure. As previous works [22, 7] have validated that the explicit-implicit hybrid structures benefit fitting speed and accuracy. We thus build a hybrid model to represent video signals. As shown in the left part of Fig. 2, our explicit module contains tri-plane and sparse grids. The former is three feature maps designed to record video features and the latter is a set of parameter grids sparsely distributed across the $x-y-t$ coordinate system of the video. Both of them are trainable parameters, denoted as $\boldsymbol{\theta}_{ex}$. Given a spatio-temporal coordinate $(x,y,t)$, we extract features from its neighbor grids and tri-plane, represented as the white cube in Fig. 2. Our implicit part is an MLP that decodes the aggregated features into a pixel value, whose parameters are $\boldsymbol{\theta}_{im}$. We update $\widetilde{\boldsymbol{\theta}}=\left\{\boldsymbol{\theta}_{im},\boldsymbol{\theta}_{ex}\right\}$ to fit a video V and then achieve editing effects.
@@
As the batch size is independent of the video length, our GPU memory demand remains constant. Notably, we supervise training with randomly sampled pixels across the entire video rather than a batch of frames, since we found this scheme yields faster convergence and slightly better fitting accuracy compared to training frame-by-frame. In this stage, NVF learns the temporal and motion priors of the given video, which facilitates subsequent consistent editing.
@@
Note that we only update NVF and fix the T2I model. Different from Sec. 3.2 that samples pixels randomly, here we supervise NVF frame-by-frame because T2I models exhibit optimal performance on complete images rather than irregularly sampled pixels. As only one frame is edited per iteration, our GPU memory requirement is constant and independent of the video length. Importantly, we adopt IP2P [4] as the T2I model, and further improve it for better editing effects, which is discussed in Sec. 3.3. It also proves that the T2I model of NVEdit is replaceable and improvable, related further experiments are given in Sec. 4.4.
In addition, a progressive optimization strategy is developed for temporal coherence. Specifically, At the beginning of the field editing stage, we assign a lower weight to the instruction condition, i.e., setting $s_{P}$ in Eq. 2 to a small number, and gradually increasing it as the optimization proceeds. This progressive strategy systematically guides the rendered frames toward the desired editing effects step-by-step, while diminishing the updating intensity to NVF at each iteration. Thus, we effectively preserve the temporal priors learned at the video fitting stage. We evaluate the contribution of this strategy in Sec. 4.4.
@@
where T is the set of all coordinates on the temporal axis. Results in Sec. 4 prove that these final rendered videos perform impressive temporal consistency. We believe that is implicitly achieved by two factors: i) In video fitting stage, NVF has learned strong temporal priors from the original video, which is inherently temporal consistent. ii) The optimization strategy in field editing stage preserves the learned temporal priors as it updates model from weak to strong smoothly.
3.3 Instruct-Pix2Pix+
Although IP2P achieves impressive image editing, for local editing, it alters the region to be retained, as shown in Fig. 3. We suppose this is an inherent flaw of IP2P as it samples based on Eq. 2, which only ensures results to be roughly consistent with the image condition, but lacks strong constraints on unedited region. Typically, users prefer instructing specific regions while leaving others untouched. Previous work [11] has used mask to restrict editable regions, but its three-step process requires redundant calculations and well-designed paired texts. To this end, we develop Instruct-Pix2Pix+ to enhance the content fidelity of edited images to the conditional image $I$, without any additional computation.
@@ 3.3 Instruct-Pix2Pix+
We observed that $\boldsymbol{\varepsilon}_{{\boldsymbol{\theta}}_{P}}$ represents the relationship between the generated image and text condition. Hence, a simple yet effective auxiliary mask is proposed. We apply normalization and dimensionality reduction to $\boldsymbol{\varepsilon}_{{\boldsymbol{\theta}}_{P}}$, getting a gray-scale image $\textbf{G}_{x,I,T}$, in which a higher value $x$ means the pixel is more likely to be relevant to instruction, and we allow it to change. In contrast, a lower $x$ signals that the pixel is not require change. A mask threshold $\tau\in[0,1]$ is set on $\textbf{G}_{x,I,T}$ to get our auxiliary mask, $\textbf{M}_{x,I,T}=\mathbbm{1}(\textbf{G}_{x,I,T}\geq\tau)$. As $\textbf{M}_{x,I,T}$ is derived from $\boldsymbol{\varepsilon}_{{\boldsymbol{\theta}}_{P}}$, a sub-term of the predicted noise in generation process. $\textbf{M}_{x,I,T}$ is regarded as an auxiliary product obtained during editing and termed as ‘auxiliary mask’.
@@
4 Experiments
4.1 Implementation Details
To generate edited frames for optimizing NVF, we use the pre-trained IP2P [4] as the base model. The proposed auxiliary mask is introduced to improve its local editing capability. Following previous works [22, 15, 36], UVG-HD [27], DAVIS [35], and some other in-the-wild videos are adopted to evaluate the editing effect.
Table 1: Quantitative evaluation of the edited results obtained from different SOTA approaches, including IP2P [4], Video-P2P [25], V2V-zero [47], TokenFlow [13], Fatezero [36], and CoDeF [33]. ‘Tem-Con’, ‘Frame-Acc’, and ‘Frame-Score’ are the objective metrics calculated by the pre-trained CLIP [38] model, while ‘Edit’, ‘Image’, and ‘Temporal’ are the subjective metrics scored by users. The best and the second-best scores are highlighted in bold and underline respectively.
@@ 4.2 Comparison with the state-of-the-art
We prove our superior effectiveness by comparing with other five State-Of-The-Art (SOTA) text-driven video editing approaches, including Video-P2P [25], Vid2Vid-zero (V2V-zero) [47], TokenFlow [13], Fatezero [36], and CoDeF [33]. Furthermore, to demonstrate that the proposed method generates more temporal consistent results than directly applying the T2I model on videos, we fix the random seed of IP2P [4] and showcase its results as a baseline. All of these methods are recently proposed. Note that for prompt-to-prompt approaches, to ensure a fair comparison, we elaborate the source-target prompt pairs for them.
Quantitative comparison. We present both the objective and subjective metrics of our method and other recent SOTA algorithms. Following [3, 36], for subjective evaluation, we also prepared 16 sets of edited results, shuffled the order of methods within each set, and asked users to rate different results in each set from 1 (the worst) to 5 (the best). As shown in Tab. 1, the proposed method achieves the best temporal consistency and frame-wise editing accuracy, and demonstrates comparable semantic alignment to the canonical image-based method [33] in the feature space of the CLIP model. As for user studies, on average, our approach obtains the best user preferences across three aspects.
@@
Figure 5: Comparison of GPU memory overhead. As frames increase, our method only adds minimal memory overhead, which allows it to edit long videos. Although CoDeF [33] consumes less GPU memory, NVEdit achieves better performance as illustrated in Tab. 1 and it is still memory-efficient compared to most existing methods.
@@
Qualitative comparison. In order to provide a more intuitive comparison, we present the visual results of NVEdit and other baselines in two different scenarios in Fig. 4, showcasing two types of motion changes: the upper scenario changes action, while the lower one changes viewpoint. Instructions used in IP2P and our method are given above. For the prompt-to-prompt methods, we designed original prompts as ‘A man is walking his dog on the road in front of the courtyard’ and ‘A person stands before the wall’, whilst the target prompts are ‘A man is walking his tiger on the road in front of the courtyard’ and ‘A person in pink shirt stands before the wall’. Observing the results, the edited results of frame-wise IP2P vary a lot among different frames (as indicated by the arrows) and fail to preserve unedited region (as shown in the bottom scene). Although other methods develop various techniques, such as attention-related control or null-text inversion, their results incorrectly modify the background and posture, or fail in certain viewpoints. Notice that CoDeF maps video content to a canonical image and edits it through ControlNet [56]. For a fair comparison, we use IP2P to edit the canonical image of CoDeF in this paper, omitting the proposed auxiliary mask as it is a novel tool invented in this work. In the top-row example, the editing of the canonical image is occasionally subpar. However, CoDeF relies entirely on this edit, leading to an unsatisfactory appearance in all of its predicted frames. In contrast, our method, with progressive optimization using numerous pseudo-GTs, accommodates errors in certain pseudo-GTs. Besides, in the lower row, IP2P incorrectly colors the wall, affecting the entire video of CoDeF, whereas our method precisely controls editable region due to IP2P+.
@@
GPU memory comparison. To demonstrate the superior GPU memory efficiency of our method, we use different algorithms to edit frames with a resolution of 480 $\times$ 480 and record their GPU memory overhead. Experiments are conducted on a single NVIDIA A800 GPU. As shown in Fig. 5, we compare different methods within 32 frames to illustrate their GPU memory growth trends. Except CoDeF [33] and our method, other approaches require inputting more frames to diffusion models for editing, resulting in a rapid increase in GPU memory. In contrast, by leveraging the efficient encoding of neural representation, both CoDeF and our method add only minor additional memory. Since CoDeF employs an implicit representation, whilst we develop an explicit-implicit hybrid structure as described in Sec. 3.2, our GPU memory overhead is higher than that of CoDeF. However, as noted in [7], hybrid model enables better representation effect than the implicit one. Although CoDeF maps video content to a canonical image, encoding only motion information, which reduces GPU memory demand. For videos with significant variations, its mapping and encoding tend to be inaccurate. While our method encodes content and motion together through an effective hybrid structure to address it. Comparative experiments also prove our SOTA performance. Our method only needs more GPU memory than CoDeF and is still memory-efficient compared to most existing methods.
Figure 6: NVEdit enables temporal consistent, high-quality edits of real-world long videos with hundreds of frames. The edited results faithfully conform to the instruction, while strictly inheriting original information, such as the posture, background, and motion.